# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v25

@@ 3.23 Sinhala Sign Language
Priyankara et al. [654] have conducted a short survey on the available Sinhala sign language recognition systems. They look at 12 studies and conclude that image processing with CNN is the most used technique for Sign Language translation. Along with a Sinhala sign language recognition system. Wijegoonaratna [655] has created a neural network-based approach for real-time Sinhala sign language gesture recognition. The approach by Hettiarachchi and Meegama [650], the approach by Dilakshan and Priyadarshana [656], as well as the approach by Peiris [648] have used CNN to recognize the fingerspelling alphabet of Sinhala sign language. Perera and Jayalal [657] also has used CNN to translate Sinhala sign Language to Sinhala text. However, they have also explicitly used Scale Invariant Feature Transform (SIFT) [658] features. A study for the limited use case of translating 15 Sinhala signs to text was conducted by Fernando and Wimalaratne [644]. The study by Charuka et al. [154] introduces a novel skeleton-based sign language recognition method named ProtoSign built upon prototypical networks [659] to do few-shot learning for Sinhala sign language recognition. The study by Haputhanthri et al. [660] used ResNet followed by a transformer encoder with multi-head attention to recognise sentence-level Sinhala sign language. The study by Fernando et al. [661] introduced a mobile app capable of translating Sinhala Sign language to text. They further enhance the results of their system with a facial emotion detection system created with CNNs. The study by Gamage et al. [662] also combines Sinhala Sign language recognition and emotion recognition. Krishnananthan et al. [663] introduced HANDTALK web-based application designed to teach sign-language. As Such it has a component that translates the sign language gestures of the learner into text. In their study Gedaragoda et al. [664] used traditional machine learning approaches to recognize Sinhala Sign Language. A pose-based method for Sinhala Sign Language recognition was introduced by Indatissa [665]. The work by Perera [666] uses CNN for Sri Lanka Sign Language recognition. The work by Nadeesha and Wasalthilaka [667] used classical machine learning techniques to recognize 40 Sinhala signs from 100 pose videos. They reported a success rate of 86.7%. Rajapakshe et al. [668] used a hybrid CNN-LSTM model and reports 94% accuracy on the 20 Sinhala signs they tested on. The work by Navodya et al. [669] reports accuracies between 87% and 92% for recognising Sinhala sign language. Thennakon et al. [670] on the other hand reports 96% accuracy in recognising real-time Sinhala Sign Language signs. The work by Kavinda et al. [671] proposed a multi-modal system that converts both sign language and miming language to Sinhala text. Further, they have a component to convert Sinhala audio to miming language and Sinhala text as well. Strides have been made in the domains of computer interpreting for written Sinhala [672] and animation of finger-spelt words and number signs [673]. The survey by Ahinsa et al. [674] discusses text-to-Sinhala sign language in the contexts of: finger-spelling and translation. It should be noted here that the translation they describe here is text-to-sign language translation and not translation between any two or more natural languages. The work by Idushan et al. [675] claims to be able to convert Sinhala audio to SSL and also recognize SSL and interpret dynamic signs.
@@
The work by Samarasinghe et al. [645] used a CNN model to convert Sinhala sign language to Sinhala text. Kumar et al. [676] introduced EasyTalk, a tool that translates Sri Lankan sign language to text as well as converts English text into Sri Lankan Sign Language. The work by dha [649] used CNN to classify features of static Sinhala signs, while dynamic signs were modelled as a tree structure of static signs. Dahanayaka et al. [647] also used CNN in their multi-model approach to recognize Sinhala sign language. Both the approaches by Perera and Kumarika [677] and Herath and Ishanka [678] integrated MediaPipe [679] with Deep Neural Networks to interpret Sri Lankan sign language. Perera and Kumarika [677] only used LSTM while Herath and Ishanka [678] used LSTM, CNN, and CNN-LSTM. Rishan et al. [642] uses Leap Motion Controller154154154https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text. Dissanayake et al. [646] have introduced Utalk, a mobile app to interpret Sri Lankan Sign Language.
@@ 3.27 Language Acquisition and Evaluation
Sandathara et al. [701] proposed a system which they named Arunalu that they claimed to use Voice recognition, Natural Language Processing, Machine Learning, and Deep Learning concepts to help individuals with dyslexia overcome problems of reading Sinhala. The learning bot MiMi proposed by Vithana et al. [702] assists children to learn to speak without stuttering. Nethmi et al. [703] introduced Narrataa, a mobile application targeted at developing the Sinhala vocabulary of children by using computer vision to caption images in Sinhala. The mobile application developed by Saranga et al. [704] helps individuals learn Sinhala letters with the help of images, colours, and shapes. The application also attempts to generate stories in Sinhala. The Katha App developed by Wijesooriya et al. [705] uses SVM and LSTm to identify phonological disorders from Sinhala speech. Francis et al. [706] claims to use Sinhala OCR and PoS tagging in their effort to build the SinhaLearn, a system for improving Sinhala Proficiency of grade 5 scholarship students. Praveen et al. [375] also claims to have created a web-based education system that helps grade 5 scholarship students who are learning the Sinhala language. They claim that they have created a Sinhala Spell checker with 82% accuracy and a Sinhala TTS system with 98% accuracy within this system. Weerasinghe et al. [707] introduced LexiLearn, a gamified system intended to help Sinhala learners with dyslexia. The survey by Perera and Sumanathilaka [708] discusses the lack of annotated datasets and NLP resources to cater for Sinhala-speaking adults with dyslexia.
@@ 3.28 Embeddings, LLMs, and Other Language Models
A two-tiered model to embed Sinhala sentences was proposed by Weeraprameshwara et al. [709]. According to their results, Seq2Seq GRU with attention run on fastText word embedding obtains the best results for Sinhala sentence embedding. Dhananjaya et al. [107] has published two Sinhala BERT models: SinBERT-large155155155https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small156156156https://huggingface.co/NLPC-UOM/SinBERT-small. A smaller BERT model named SinhalaBERTo157157157https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper. This BERT model has been created by training vanilla RoBERTa [710] on the OSCAR [155, 156] dataset. Demotte et al. [276] have released the word embeddings158158158https://drive.google.com/drive/folders/1CgqPIloW5DKLVj3WViMitIhtrswOv_qR as well as the code159159159https://github.com/theisuru/sentiment-tagger for their word embedding-based Sinhala sentiment analyzer. Gurgurov et al. [711] included Sinhala as one of their 87 low-resource languages for which they provided word embeddings in their LowREm repository. They credit Demotte et al. [276] as their source for Sinhala embeddings. Gurgurov et al. [712]160160160https://github.com/d-gurgurov/Knowledge-Driven-Adaptation-LLMs argued that mLMs such as mBERT and XLM-R are better suited for low-resourced languages such as Sinhala as opposed to LLMs and suggested parameter-efficient adapter-based methods for adapting such mLMs.
@@
Wickramasinghe and de Silva [148] used Sinhala-English dictionary data [147] for obtaining Sinhala-English embedding alignment. Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese. The relevant code and data can be accessed on GitHub161161161https://bit.ly/3t3SKu7. Subsequent work by Rathnayake et al. [713] also attempts BLI for Sinhala English but uses UVecMap (Unsupervised VecMap [714]) as the baseline. They report that UVecMap combined with Linear Transformation archives the best result of 33.18 for En-Si Word2Vec embedding, while UVecMap combined with Linear Transformation and CSCBLI [715] archives the best result of 29.49 for En-Si FastText embedding.
@@
Üstün et al. [171] introduced Aya, a massively multilingual generative language model162162162https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [159]. Jayakody and Dias [716] conducted a comparative analysis on LLMs for Sinhala. In it, they observed that the small (7̃B) models of Claude163163163https://docs.anthropic.com/en/docs/about-claude/models and GPT-4o164164164https://openai.com/index/hello-gpt-4o/ handled Sinhala better than Llama 3 (8B)165165165https://llama.meta.com/docs/get-started/ and Mistral 7B166166166https://docs.mistral.ai/. Further, they reported that they had to exclude Falcon 2167167167https://falconllm.tii.ae/ and Gemini168168168https://deepmind.google/technologies/gemini/ from the study due to poor or non-existent support for Sinhala. Hansaka and Sumanathilaka [717] also claims to have conducted a review on fine-tuning approaches for LLMs such as GPT-3, GPT-4, and LlaMA in the context of Sinhala. Chang et al. [718] introduced Goldfish, a collection of monolingual language models169169169https://huggingface.co/goldfish-models which includes Sinhala trained on data from MADLAD [160], NLLB [134], and Glot500 [109]. The work by Ayesha et al. [719] used a fined-tuned Llama 3.1 (8B) model [720] to predict student performance by analyzing textual entries created by students in Sinhala and English. Tang et al. [721] evaluated the susceptibility of LLMs being tricked into answering illegal questions in low-resource languages when the same questions would have been successfully rejected had they been asked in a high-resourced language. They provide the framework for testing with the results on eight low-resourced languages, which include Sinhala.
@@
Vayani et al. [164] analyzed the performance of a number of LLMs over 100 languages that included Sinhala. The basic dataset was created by machine translating English QA pairs from LLaVA-Bench (In-the-Wild) dataset [165]. This was further extended with culturally specific data points by searching for images under country name, language name, and cultural category. This dataset, which they named ALM-Bench170170170https://huggingface.co/datasets/MBZUAI/ALM-Bench, is publicly available. They made a number of observations in this aspect. First, they observed that for the Q&A task, the performance of GPT-4o [166] for Sinhala is better than Qwen2-VL [722]. Next, they report that including images in the question further improved Sinhala results of GPT-4o by 38.7%. In the error analysis of the same, they report that 36% of the errors are language errors (fluency), 31% of the errors are due to lack of knowledge, and 18% of the errors are due to a lack of cultural understanding.
@@ 3.29 Text Simplification
Ranathunga et al. [723] introduced SiTSE, a Sinhala text simplification dataset and the associated benchmark. They claim that the data set is publicly available on github171171171https://github.com/brainsharks-fyp17/Sinhala-Text-Simplification-Dataset-and-Evaluation; however, currently it resolves to a 404. The text simplification Shared task of Building Educational Applications (BEA) at MLSP 2024 [724] included the NSINA [108] Sinhala corpus. The aggregated data from the shared task is available online as Train-Test data172172172https://github.com/MLSP2024/MLSP_Data/ and cleaned gold data173173173https://huggingface.co/datasets/MLSP2024/MLSP2024. Enomoto et al. [725] experimented with GPT-4 and noted that re-ranking systems outperformed the base system for Sinhala while the opposite was true for all the other nine languages. Goswami et al. [726] also tested GPT-4 (Terbo) for this task against the top 10 suggestion model and reported better results for GPT-4 for Sinhala than the top 10 suggestion model. Seneviratne and Suominen [727] compared multiple models for the English-Sinhala language pair simplification for the lexical complexity prediction task and lexical substitution task. In the submission by Cristea and Nisioi [728] word frequency and other meta-knowledge are not used for Sinhala text simplification while they are used for the other languages in the task. Their Sihala models are trained only using Sinhala text Google translated to English. As such the results they report for Sinhala for this task is quite weak compared to their results for the other languages.

### v26

@@ 3.23 Sinhala Sign Language
Priyankara et al. [698] have conducted a short survey on the available Sinhala sign language recognition systems. They look at 12 studies and conclude that image processing with CNN is the most used technique for Sign Language translation. Along with a Sinhala sign language recognition system. Wijegoonaratna [699] has created a neural network-based approach for real-time Sinhala sign language gesture recognition. The approach by Hettiarachchi and Meegama [694], the approach by Dilakshan and Priyadarshana [700], as well as the approach by Peiris [692] have used CNN to recognize the fingerspelling alphabet of Sinhala sign language. Perera and Jayalal [701] also has used CNN to translate Sinhala sign Language to Sinhala text. However, they have also explicitly used Scale Invariant Feature Transform (SIFT) [702] features. A study for the limited use case of translating 15 Sinhala signs to text was conducted by Fernando and Wimalaratne [688]. The study by Charuka et al. [157] introduces a novel skeleton-based sign language recognition method named ProtoSign built upon prototypical networks [703] to do few-shot learning for Sinhala sign language recognition. The study by Haputhanthri et al. [704] used ResNet followed by a transformer encoder with multi-head attention to recognise sentence-level Sinhala sign language. The study by Fernando et al. [705] introduced a mobile app capable of translating Sinhala Sign language to text. They further enhance the results of their system with a facial emotion detection system created with CNNs. The study by Gamage et al. [706] also combines Sinhala Sign language recognition and emotion recognition. Krishnananthan et al. [707] introduced HANDTALK web-based application designed to teach sign-language. As Such it has a component that translates the sign language gestures of the learner into text. In their study Gedaragoda et al. [708] used traditional machine learning approaches to recognize Sinhala Sign Language. A pose-based method for Sinhala Sign Language recognition was introduced by Indatissa [709]. The work by Perera [710] uses CNN for Sri Lanka Sign Language recognition. The work by Nadeesha and Wasalthilaka [711] used classical machine learning techniques to recognize 40 Sinhala signs from 100 pose videos. They reported a success rate of 86.7%. Rajapakshe et al. [712] used a hybrid CNN-LSTM model and reports 94% accuracy on the 20 Sinhala signs they tested on. The work by Navodya et al. [713] reports accuracies between 87% and 92% for recognising Sinhala sign language. Thennakon et al. [714] on the other hand reports 96% accuracy in recognising real-time Sinhala Sign Language signs. The work by Kavinda et al. [715] proposed a multi-modal system that converts both sign language and miming language to Sinhala text. Further, they have a component to convert Sinhala audio to miming language and Sinhala text as well. Strides have been made in the domains of computer interpreting for written Sinhala [716] and animation of finger-spelt words and number signs [717]. The survey by Ahinsa et al. [718] discusses text-to-Sinhala sign language in the contexts of: finger-spelling and translation. It should be noted here that the translation they describe here is text-to-sign language translation and not translation between any two or more natural languages. Thrimahavithana et al. [719] used 3D avatars to show Sinhala sign language on screen. Their system was evaluated with a usability score, error rate in text-to-sign language conversion, and overall user satisfaction. The work by Idushan et al. [720] claims to be able to convert Sinhala audio to SSL and also recognize SSL and interpret dynamic signs.
@@
The work by Samarasinghe et al. [689] used a CNN model to convert Sinhala sign language to Sinhala text. Kumar et al. [721] introduced EasyTalk, a tool that translates Sri Lankan sign language to text as well as converts English text into Sri Lankan Sign Language. The work by dha [693] used CNN to classify features of static Sinhala signs, while dynamic signs were modelled as a tree structure of static signs. Dahanayaka et al. [691] also used CNN in their multi-model approach to recognize Sinhala sign language. Both the approaches by Perera and Kumarika [722] and Herath and Ishanka [723] integrated MediaPipe [724] with Deep Neural Networks to interpret Sri Lankan sign language. Perera and Kumarika [722] only used LSTM while Herath and Ishanka [723] used LSTM, CNN, and CNN-LSTM. Rishan et al. [686] uses Leap Motion Controller194194194https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text. Dissanayake et al. [690] have introduced Utalk, a mobile app to interpret Sri Lankan Sign Language. Jayasekara et al. [725] uses EfficientNet-B0 model [726] to detect static Sri Lankan Sign Language.
@@ 3.27 Language Acquisition and Evaluation
Sandathara et al. [748] proposed a system which they named Arunalu that they claimed to use Voice recognition, Natural Language Processing, Machine Learning, and Deep Learning concepts to help individuals with dyslexia overcome problems of reading Sinhala. The learning bot MiMi proposed by Vithana et al. [749] assists children to learn to speak without stuttering. Nethmi et al. [750] introduced Narrataa, a mobile application targeted at developing the Sinhala vocabulary of children by using computer vision to caption images in Sinhala. The mobile application developed by Saranga et al. [751] helps individuals learn Sinhala letters with the help of images, colours, and shapes. The application also attempts to generate stories in Sinhala. The Katha App developed by Wijesooriya et al. [752] uses SVM and LSTm to identify phonological disorders from Sinhala speech. Francis et al. [753] claims to use Sinhala OCR and PoS tagging in their effort to build the SinhaLearn, a system for improving Sinhala Proficiency of grade 5 scholarship students. Praveen et al. [393] also claims to have created a web-based education system that helps grade 5 scholarship students who are learning the Sinhala language. They claim that they have created a Sinhala Spell checker with 82% accuracy and a Sinhala TTS system with 98% accuracy within this system. Weerasinghe et al. [754] introduced LexiLearn, a gamified system intended to help Sinhala learners with dyslexia. The survey by Perera and Sumanathilaka [755] discusses the lack of annotated datasets and NLP resources to cater for Sinhala-speaking adults with dyslexia. The work by Perera and Sumanathilaka [756] created a pipeline with: Whisper for speech-to-text, SinBERT for dyslexic error detection, mT5 and Mistral to generate corrected text, and finally GoogleTTS (gTTS) for text-to-speech.
@@ 3.28 Embeddings, LLMs, and Other Language Models
A two-tiered model to embed Sinhala sentences was proposed by Weeraprameshwara et al. [757]. According to their results, Seq2Seq GRU with attention run on fastText word embedding obtains the best results for Sinhala sentence embedding. Dhananjaya et al. [110] has published two Sinhala BERT models: SinBERT-large195195195https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small196196196https://huggingface.co/NLPC-UOM/SinBERT-small. A smaller BERT model named SinhalaBERTo197197197https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper. This BERT model has been created by training vanilla RoBERTa [758] on the OSCAR [158, 159] dataset. Demotte et al. [289] have released the word embeddings198198198https://drive.google.com/drive/folders/1CgqPIloW5DKLVj3WViMitIhtrswOv_qR as well as the code199199199https://github.com/theisuru/sentiment-tagger for their word embedding-based Sinhala sentiment analyzer. Gurgurov et al. [759] included Sinhala as one of their 87 low-resource languages for which they provided word embeddings in their LowREm repository. They credit Demotte et al. [289] as their source for Sinhala embeddings. Gurgurov et al. [760]200200200https://github.com/d-gurgurov/Knowledge-Driven-Adaptation-LLMs argued that mLMs such as mBERT and XLM-R are better suited for low-resourced languages such as Sinhala as opposed to LLMs and suggested parameter-efficient adapter-based methods for adapting such mLMs. Enevoldsen et al. [761] introduced MMTEB201201201https://github.com/embeddings-benchmark/mteb, a multilingual embedding benchmark along with a leader board202202202https://huggingface.co/spaces/mteb/leaderboard. They used the news classification data set created by de Silva [30] for Sinhala. Ranasinghe et al. [762] introduced SINHALA-GLUE203203203https://github.com/Sinhala-NLP/Sinhala-GLUE, a language understanding benchmark for Sinhala, along with three new trained Sinhala embedding models: Sinhala-BERT (Raja), Sinhala-RoBERTa (Koliya), and Sinhala-Electra (Mahasen). These models are not publicly available.
@@
Wickramasinghe and de Silva [151] used Sinhala-English dictionary data [150] for obtaining Sinhala-English embedding alignment. Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese. The relevant code and data can be accessed on GitHub204204204https://bit.ly/3t3SKu7. Subsequent work by Rathnayake et al. [763] also attempts BLI for Sinhala English but uses UVecMap (Unsupervised VecMap [764]) as the baseline. They report that UVecMap combined with Linear Transformation archives the best result of 33.18 for En-Si Word2Vec embedding, while UVecMap combined with Linear Transformation and CSCBLI [765] archives the best result of 29.49 for En-Si FastText embedding. The following work by Wickramasinghe and de Silva [766] compared multiple BLI techniques across different language pairs and showed that LaBSE [767] with cross-domain similarity local scaling (CSLS) [768] generally obtained the best result. Further, they proposed Vocabulary Pruning and stemming when one or both of the languages considered is highly inflected as is the case with Sinhala. Yu [769] uses FLORES-200 [137] data set to show that their cross-model semantic alignment method is superior to CLIP [770] and ALIGN [771]. They specifically point out Sinhala as one of the languages that shows the most improvement for recall@1 and Mean Rank.
@@
Pramodya et al. [772] released, SinhalaMMLU, a data set of more than 7000 Sinhala questions and answers covering secondary to collegiate education levels in Sri Lanaka. They used this dataset to evaluate 26 LLMs and reported that Claude 3.5 achieved the highest average accuracy of 67%. Üstün et al. [173] introduced Aya, a massively multilingual generative language model205205205https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [162]. Amarakoon and Rajapakse [773] further fine-tuned the Aya model using Sri Lankan business and regulatory datasets to automate multilingual customer interactions and generate context-sensitive business insights in Sinhala and Tamil for Sri Lankan Small and Medium Enterprises (SMEs). Aravinda et al. [774] introduced SinLlama206206206https://huggingface.co/polyglots/SinLlama_v01, the first pre-trained Sinhala LLM-base model. They used continual pre-training with a 10-million-sentence Sinhala corpus on Llama-3-8B to obtain this model. They show the utility of this model with three text classification tasks. Further, for this task, they created a Sinhala tokeniser, which is also publicly available207207207https://huggingface.co/polyglots/Extended-Sinhala-LLaMA. Jayakody and Dias [775] conducted a comparative analysis on LLMs for Sinhala. In it, they observed that the small (7̃B) models of Claude208208208https://docs.anthropic.com/en/docs/about-claude/models and GPT-4o209209209https://openai.com/index/hello-gpt-4o/ handled Sinhala better than Llama 3 (8B)210210210https://llama.meta.com/docs/get-started/ and Mistral 7B211211211https://docs.mistral.ai/. Further, they reported that they had to exclude Falcon 2212212212https://falconllm.tii.ae/ and Gemini213213213https://deepmind.google/technologies/gemini/ from the study due to poor or non-existent support for Sinhala. Hansaka and Sumanathilaka [776] also claims to have conducted a review on fine-tuning approaches for LLMs such as GPT-3, GPT-4, and LlaMA in the context of Sinhala. Chang et al. [777] introduced Goldfish, a collection of monolingual language models214214214https://huggingface.co/goldfish-models which includes Sinhala trained on data from MADLAD [163], NLLB [137], and Glot500 [112]. The work by Ayesha et al. [778] used a fined-tuned Llama 3.1 (8B) model [779] to predict student performance by analyzing textual entries created by students in Sinhala and English. Tang et al. [780] evaluated the susceptibility of LLMs being tricked into answering illegal questions in low-resource languages when the same questions would have been successfully rejected had they been asked in a high-resourced language. They provide the framework for testing with the results on eight low-resourced languages, which include Sinhala. In their work to overcome data sparsity for LLMs, Lai et al. [781]215215215Later elaborated as Lai [782]. tested the Sinhala text generation and understanding results of Google Translate against LLaMA, BLOOM, and several other variations. Even though they point to a repository of the instruction data set they created216216216https://github.com/boschresearch/ACL24-MLLM, currently, the data set cannot be found at the given URL. Puranegedara et al. [783] proposed an intermediate layer fusing method to LLMs so that they would perform better for low-resource languages such as Sinhala. They show their improvement with Sinhala News Results on the data set created by de Silva [30].
@@
Vayani et al. [167] analyzed the performance of a number of LLMs over 100 languages that included Sinhala. The basic dataset was created by machine translating English QA pairs from LLaVA-Bench (In-the-Wild) dataset [168]. This was further extended with culturally specific data points by searching for images under country name, language name, and cultural category. This dataset, which they named ALM-Bench217217217https://huggingface.co/datasets/MBZUAI/ALM-Bench, is publicly available. They made a number of observations in this aspect. First, they observed that for the Q&A task, the performance of GPT-4o [169] for Sinhala is better than Qwen2-VL [784]. Next, they report that including images in the question further improved Sinhala results of GPT-4o by 38.7%. In the error analysis of the same, they report that 36% of the errors are language errors (fluency), 31% of the errors are due to lack of knowledge, and 18% of the errors are due to a lack of cultural understanding. Shafique et al. [785] when introducing ViMUL-Bench, a multimodal video Q&A benchmark for LLMs that includes Sinhala, show that GPT4o has the best results for Sinhala.
@@ 3.29 Text Simplification
The MultiLS data set218218218https://huggingface.co/datasets/MLSP2024/MLSP2024 introduced by Shardlow et al. [786] for MLSP 2024 Shared Task219219219https://github.com/MLSP2024/MLSP_Data includes a Sinhala text simplification data set with data from News and Religious domains. Ranathunga et al. [787] introduced SiTSE, a Sinhala text simplification dataset which is publicly available on github220220220https://github.com/brainsharks-fyp17/Sinhala-Text-Simplification-Dataset-and-Evaluation. It has 1000 complex Sinhala texts along with two separate simplifications for each of them. Further, they provide the evaluation metrics that can be used to evaluate any subsequent research conducted using the same data.
The text simplification Shared task of Building Educational Applications (BEA) at MLSP 2024 [788] included the NSINA [111] Sinhala corpus. The aggregated data from the shared task is available online as Train-Test data221221221https://github.com/MLSP2024/MLSP_Data/ and cleaned gold data222222222https://huggingface.co/datasets/MLSP2024/MLSP2024. Enomoto et al. [789] experimented with GPT-4 and noted that re-ranking systems outperformed the base system for Sinhala while the opposite was true for all the other nine languages. Goswami et al. [790] also tested GPT-4 (Terbo) for this task against the top 10 suggestion model and reported better results for GPT-4 for Sinhala than the top 10 suggestion model. Seneviratne and Suominen [791] compared multiple models for the English-Sinhala language pair simplification for the lexical complexity prediction task and lexical substitution task. In the submission by Cristea and Nisioi [792] word frequency and other meta-knowledge are not used for Sinhala text simplification while they are used for the other languages in the task. Their Sihala models are trained only using Sinhala text Google translated to English. As such the results they report for Sinhala for this task is quite weak compared to their results for the other languages. Even though the main focus on the work by Nohejl et al. [793] is introducing a Japanese text simplification data set and a model, to demonstrate the generalizability of the system, they use the MultiLS data set [786] which includes Sinhala.